OpenStudy (chosenmatt):

what is 4 1/5 + 3 1/10

OpenStudy (anonymous):

4 1/5 = 21/5 = 42/10 3 1/10 = 31/10 so 42/10 + 31/10 = ?

OpenStudy (anonymous):

7 3/10

OpenStudy (midhun.madhu1987):

 take the LCM and simplify